2. Two smallcharged spheres are placed in vacuum on the x-axis: +3pCatx-Oand-5,Cat x-40cm, where must a third charge q be placed if the force it experiences is to be zero?
Answer #1

Let us consider that the 3rd particle is placed left to the +3uC at a distance d

Now in order to get zero force, net electric at that point must be zero i.e. Electric field due to +3 uC is equal and opposite to - 5uC

k q1 / d^2 = k q2 / ( d + 40)^2

1 + 40/d = sqrt ( q2 / q1 )

1 + 40/d = sqrt ( 5/3)

d = 137.46 cm

X co ordinate of this point

X = - 137.46 cm

=========

Do comment in case any doubt, will reply for sure . Goodluck